# Two Heads are Better than One: Robust Learning Meets Multi-branch Models

### v3

@@ Abstract
Deep neural networks (DNNs) are vulnerable to adversarial examples, in which DNNs are misled to false outputs due to inputs containing imperceptible perturbations. Adversarial training, a reliable and effective method of defense, may significantly reduce the vulnerability of neural networks and becomes the de facto standard for robust learning. While many recent works practice the data-centric philosophy, such as how to generate better adversarial examples or use generative models to produce additional training data, we look back to the models themselves and revisit the adversarial robustness from the perspective of deep feature distribution as an insightful complementarity. In this paper, we propose Branch Orthogonality adveRsarial Training (BORT) to obtain state-of-the-art performance with solely the original dataset for adversarial training. To practice our design idea of integrating multiple orthogonal solution spaces, we leverage a simple and straightforward multi-branch neural network that eclipses adversarial attacks with no increase in inference time. We heuristically propose a corresponding loss function, branch-orthogonal loss, to make each solution space of the multi-branch model orthogonal. We evaluate our approach on CIFAR-10, CIFAR-100 and SVHN against $\ell_{\infty}$ norm-bounded perturbations of size $\epsilon=8/255$, respectively. Exhaustive experiments are conducted to show that our method goes beyond all state-of-the-art methods without any tricks. Compared to all methods that do not use additional data for training, our models achieve 67.3% and 41.5% robust accuracy on CIFAR-10 and CIFAR-100 (improving upon the state-of-the-art by +7.23% and +9.07%). We also outperform methods using a training set with a far larger scale than ours.
@@ I Introduction
Deep Neural Networks (DNNs) have got great research attention and success in various application scenarios [1, 2, 3, 4, 5, 6, 7, 8]. However, it is found that DNNs are vulnerable to adversarial examples [9, 10, 11, 12, 13]. Leaving the model to produce the wrong output as the attacker wishes poses a significant security risk. Therefore, the vulnerability of DNNs to adversarial attacks is a critical topic in many areas with safety-vital nature, such as autonomous driving [14, 15, 16], and medical diagnostics [17, 18, 19].
@@
Figure 2: Robust accuracy of models against AutoAttack[20] on CIFAR10 (left) and CIFAR100 (right) dataset.
Ever since Szegedy et al. [21] brought this problem to the forefront, many studies have sprung up on generating adversarial examples [22, 23] and building defenses against these attacks [24, 25]. Among many successful defenses, adversarial training became one of the mainstream and proved to be effective and practical. Madry et al. [24] first introduced projected gradient descent (PGD) to generate adversarial examples for training, achieving model robustness and invariance to input perturbations. As a standard for training adversarially robust models, a variety of modifications over their implementation have been proposed, such as Lipschitz regularization [26, 27], computationally efficient adversarial training [28, 29], utilization of unlabeled data [30, 31] and better trade-offs between robustness and natural generalization [32, 33].
However, in recent years, the improvement in robustness of the model seems to be less effective and practical. To further advance the model robustness, Carmon et al. [30], Najafi et al. [34] and Zhai et al. [35] pioneered the use of additional data (under both supervised and semi-supervised settings). Gowal et al. [36] trained a model on CIFAR-10 against $\ell_{\infty}$ perturbations of size $\epsilon=8/255$ with a maximum enhancement of $8.73\%$ robust accuracy when using additional unlabeled data, uncovering the importance of additional datasets for robustness of the model obtained from adversarial training.
@@ II-A Adversarial Attacks
Research on DNN robustness primarily involves adversarial attacks and defenses. Adversarial attacks generate subtle perturbations to input data, causing models to misclassify. These attacks are categorized as white-box (with full model access) or black-box (without model access), and as targeted (forcing a specific misclassification) or untargeted (any misclassification). Goodfellow, Shlens, and Szegedy [9] proposed the Fast Gradient Sign Method (FGSM) which generates adversarial examples. It was followed by R+FGSM [37], which adds a randomization step, and the Basic Iterative Method (BIM) [38], which takes multiple smaller gradient steps. In this paper, we mainly focus on AutoAttack [20], a widely used robustness benchmark that combines four complementary attacks, to evaluate and compare our method.
@@ II-B Adversarial Training as a defense
In response to the rapid development of adversarial attacks, researchers have proposed various defense methods, mainly categorized as certified and empirical defenses. Certified defenses use optimized objectives to guarantee robustness [39, 40, 41], while empirical defenses train models with adversarial samples.
@@
PGD [24] is a widely used method for generating adversarial perturbations by iteratively updating inputs and projecting them onto an $l_{p}$-norm ball, making it more effective than single-step attacks like FGSM [9].
Beyond adversarial training, various non-certified defenses (e.g., distillation [42], activation pruning [43], input transformation [44]) have been proposed, but many suffer from obfuscated gradients[45] and can be bypassed by advanced attacks. In contrast, adversarial training methods such as PAT[24], TRADES [32], and MART [46] remain effective against both white-box and black-box attacks. Recent work by Gowal et al.[36] revealed a robustness gap between models trained with and without extra data like 80M-TI. Follow-up studies[47, 48] showed that even synthetic data from generative models and various augmentation strategies can further enhance robustness.
@@ III Proposed Method
Notation. $(x,y)\in\mathcal{X}$ is the natural dataset, and $x^{\prime}$ signifies perturbed examples. $f^{k}$ denotes any of the $K$ branches and $f_{i}^{k}$ especially stands for the i-th ‘block’ (consistent to the definition in Zagoruyko and Komodakis [49]) of branch $k$. $\delta$ indicates perturbations bounded in the $\epsilon$-ball, where $\epsilon$ represents the numerical limit of the perturbation.
@@ III-C Optimization Formulation
Loss for outer minimization. The outer minimization problem is to allow the model to correctly classify adversarial examples. To obtain a better trade-off between the model robustness and its accuracy on clean data, we use the loss design proposed in TRADES[32]. In addition to the branch-orthogonal loss we introduced for orthogonality learning, the loss is composed of two parts: the first encourages the natural error (on clean data) to be minimized by applying the cross-entropy loss between the prediction of clean example $f^{k}(x)$ and labels $y$, and the second acts like a regularization term optimizing the difference between the prediction of clean example $f(x)$ and that of adversarial example $f^{k}(x^{\prime})$. And there is a hyper-parameter balancing the importance of the two parts.
@@ IV-A Evaluation Setting
CIFAR-10 Wu, Xia, and Wang[50] WRN-28-10(RST-AWP) 88.3% 60.0%
Gowal et al.[36] WRN-70-16 85.3% 57.2%
Rebuffi et al.[51] WRN-28-10 86.1% 57.5%
Rebuffi et al.[51] WRN-34-10 86.2% 58.1%
Rebuffi et al.[51] WRN-70-16 87.3% 60.1%
Sehwag et al.[52] + 10M DDPM WRN-34-10 86.7% 60.3%
Pang et al.[53] + 1M DDPM WRN-28-10 88.6% 61.0%
Pang et al.[53] + 1M DDPM WRN-70-16 89.0% 63.4%
Jia et al.[54] WRN-34-10(LAS-AWP) 87.7% 55.5%
Jia et al.[54] WRN-70-16 85.7% 57.6%
@@
CIFAR-100 Wu, Xia, and Wang[50] WRN-34-10 60.4% 28.9%
Gowal et al.[36] WRN-70-16 60.9% 30.0%
Rebuffi et al.[51] WRN-28-10 63.0% 29.8%
Rebuffi et al.[51] WRN-70-16 65.8% 32.4%
Sehwag et al.[52] + 1M DDPM WRN-34-10 65.9% 31.2%
Pang et al.[53] + 1M DDPM WRN-28-10 63.7% 31.1%
Pang et al.[53] + 1M DDPM WRN-70-16 65.6% 33.1%
Jia et al.[54] WRN-34-10 64.9% 30.8%
@@
SVHN Wu, Xia, and Wang[50] WRN-34-10 92.8% 58.7%
Gowal et al.[36] WRN-28-10 92.9% 57.3%
Gowal et al.[36] WRN-34-10 93.1% 58.0%
Rebuffi et al.[51] WRN-28-10 94.5% 57.3%
@@
FGSM, PGD, and AutoAttack are utilized to evaluate our proposed method along with state-of-the-arts. Testing robustness against PGD is done in 40/100 iterations with the maximum $\epsilon=8/255$ and step sizes of $0.01$ for each iteration. FGSM attack also uses the same 8/255 bound for perturbation. We leverage the implementation of AutoAttack from torchattacks repository [55].
Model training details. We use WideResNets as our base models: WRN-28-10, WRN-34-10 and WRN-70-16 [49], respectively. Each branch is trained for 50 epochs using stochastic gradient descent (SGD) as the optimizer, with an initial learning rate of 0.1 that decays to 0.01 at the 40th epoch. The hyper-parameters are configured as $\lambda_{1}=\lambda_{2}=3$ during the first 40 epochs, followed by $\lambda_{1}=\lambda_{2}=1$ for the remaining 10 epochs.
@@
TABLE III: Ablation studies of loss terms. Vanilla: Models trained entirely with clean samples. TRADES-4: The same model structure as our method proposed, but the branches are trained according to TRADES[32](differ in loss function).
@@ IV-B Overall Performance
To illustrate the effectiveness of the proposed BORT, we train various models using this framework and evaluate their robustness against different adversarial attack algorithms. Furthermore, the proposed method is compared with different state-of-the-art approaches [50, 36, 54, 52, 53, 51]. The detailed experiment results are illustrated in Table I. For a fair comparison, only the results that do not use the additional dataset (80M-Ti for example) are presented in the table, although we still outperform the results using it. The 80M-Ti dataset has been taken down by the author team, given that it may contain offensive and biased content. Likewise, we do not try to train stronger models utilizing this dataset as additional training data.
@@ V Conclusion
In this paper, we present BORT, the first adversarial training framework leveraging a multi-branch neural network, which improves network representation through multiple orthogonal branches while keeping inference time and computational cost from increasing, greatly improving model robustness against adversarial attacks. Through comprehensive experiments, we first illustrate that our proposed method goes well beyond the current state-of-the-art by a noticeable gap without the need for additional training datasets. The efficiency of multi-branch structure and branch-orthogonal loss for enhancing overall robustness was then discussed. Our method achieves top-rank performance on RobustBench under AutoAttack.

### v4

@@ Abstract
Deep neural networks (DNNs) are vulnerable to adversarial examples, in which DNNs are misled to false outputs due to inputs containing imperceptible perturbations. Adversarial training, a reliable and effective method of defense, may significantly reduce the vulnerability of neural networks and becomes the de facto standard for robust learning. While many recent works practice the data-centric philosophy, such as how to generate better adversarial examples or use generative models to produce additional training data, we look back to the models themselves and revisit the adversarial robustness from the perspective of deep feature distribution as an insightful complementarity. In this paper, we propose Branch Orthogonality adveRsarial Training (BORT) to obtain state-of-the-art performance with solely the original dataset for adversarial training. To practice our design idea of integrating multiple orthogonal solution spaces, we leverage a simple multi-branch neural network and propose a corresponding loss function, branch-orthogonal loss, to make each solution space of the multi-branch model orthogonal. We evaluate our approach on CIFAR-10, CIFAR-100 and SVHN against $\ell_{\infty}$ norm-bounded perturbations of size $\epsilon=8/255$, respectively. Exhaustive experiments are conducted to show that our method goes beyond all state-of-the-art methods without any tricks. Compared to all methods that do not use additional data for training, our models achieve 67.3% and 41.5% robust accuracy on CIFAR-10 and CIFAR-100 (improving upon the state-of-the-art by +7.23% and +9.07%).
@@ I Introduction
Deep Neural Networks (DNNs) have got great research attention and success in various application scenarios [lin2024efficient, yuan2025constructing, peng2024sums, lin2024adaptsfl, fang2024automated, lin2025hierarchical, duan2025sample, hu2024accelerating]. However, it is found that DNNs are vulnerable to adversarial examples [zhang2025robust, duan2025rethinking, du2025dp, zhang2025state]. Leaving the model to produce the wrong output as the attacker wishes poses a significant security risk. Therefore, the vulnerability of DNNs to adversarial attacks is a critical topic in many areas with safety-vital nature, such as autonomous driving [wang2021interpreting, fang2024ic3m, lin2025hsplitlora], and medical diagnostics [de2018clinically, fang2025dynamic, tang2024merit].
@@
Figure 2: Robust accuracy of models against AutoAttack[croce2020reliable] on CIFAR10 (left) and CIFAR100 (right) dataset.
Ever since Szegedy et al. [szegedy2013intriguing] brought this problem to the forefront, many studies have sprung up on generating adversarial examples [athalye2018synthesizing] and building defenses against these attacks [jeddi2020learn2perturb]. Among many successful defenses, adversarial training became one of the mainstream and proved to be effective and practical. Madry et al. [madry2017towards] first introduced projected gradient descent (PGD) to generate adversarial examples for training, achieving model robustness and invariance to input perturbations. As a standard for training adversarially robust models, a variety of modifications over their implementation have been proposed, such as Lipschitz regularization [yan2018deep], computationally efficient adversarial training [wong2020fast], utilization of unlabeled data [carmon2019unlabeled] and better trade-offs between robustness and natural generalization [zhang2020attacks].
However, in recent years, the improvement in robustness of the model seems to be less effective and practical. To further advance the model robustness, Carmon et al. [carmon2019unlabeled], Najafi et al. [najafi2019robustness] and Zhai et al. [zhai2019adversarially] pioneered the use of additional data (under both supervised and semi-supervised settings). Gowal et al. [gowal2020uncovering] trained a model on CIFAR-10 against $\ell_{\infty}$ perturbations of size $\epsilon=8/255$ with a maximum enhancement of $8.73\%$ robust accuracy when using additional unlabeled data, uncovering the importance of additional datasets for robustness of the model obtained from adversarial training.
@@ II-A Adversarial Attacks
Research on DNN robustness primarily involves adversarial attacks and defenses. Adversarial attacks generate subtle perturbations to input data, causing models to misclassify. These attacks are categorized as white-box (with full model access) or black-box (without model access), and as targeted (forcing a specific misclassification) or untargeted (any misclassification). Goodfellow, Shlens, and Szegedy [goodfellow2014explaining] proposed the Fast Gradient Sign Method (FGSM) which generates adversarial examples. It was followed by R+FGSM [tramer2017ensemble], which adds a randomization step, and the Basic Iterative Method (BIM) [kurakin2018adversarial], which takes multiple smaller gradient steps. In this paper, we mainly focus on AutoAttack [croce2020reliable], a widely used robustness benchmark that combines four complementary attacks, to evaluate and compare our method.
@@ II-B Adversarial Training as a defense
In response to the rapid development of adversarial attacks, researchers have proposed various defense methods, mainly categorized as certified and empirical defenses. Certified defenses use optimized objectives to guarantee robustness [hein2017formal, raghunathan2018certified, sinha2017certifiable], while empirical defenses train models with adversarial samples.
@@
PGD [madry2017towards] is a widely used method for generating adversarial perturbations by iteratively updating inputs and projecting them onto an $l_{p}$-norm ball, making it more effective than single-step attacks like FGSM [goodfellow2014explaining].
Beyond adversarial training, various non-certified defenses (e.g., distillation [papernot2016distillation], activation pruning [dhillon2018stochastic], input transformation [guo2017countering]) have been proposed, but many suffer from obfuscated gradients[athalye2018obfuscated] and can be bypassed by advanced attacks. In contrast, adversarial training methods such as PAT[madry2017towards], TRADES [zhang2019theoretically], and MART [wang2019improving] remain effective against both white-box and black-box attacks. Recent work by Gowal et al.[gowal2020uncovering] revealed a robustness gap between models trained with and without extra data like 80M-TI. Follow-up studies[gowal2021improving, rebuffi2021fixing] showed that even synthetic data from generative models and various augmentation strategies can further enhance robustness.
@@ III Proposed Method
Notation. $(x,y)\in\mathcal{X}$ is the natural dataset, and $x^{\prime}$ signifies perturbed examples. $f^{k}$ denotes any of the $K$ branches and $f_{i}^{k}$ especially stands for the i-th ‘block’ of branch $k$. $\delta$ indicates perturbations bounded in the $\epsilon$-ball, where $\epsilon$ represents the numerical limit of the perturbation.
@@ III-C Optimization Formulation
Loss for outer minimization. The outer minimization problem is to allow the model to correctly classify adversarial examples. To obtain a better trade-off between the model robustness and its accuracy on clean data, we use the loss design proposed in TRADES[zhang2019theoretically]. In addition to the branch-orthogonal loss we introduced for orthogonality learning, the loss is composed of two parts: the first encourages the natural error (on clean data) to be minimized by applying the cross-entropy loss between the prediction of clean example $f^{k}(x)$ and labels $y$, and the second acts like a regularization term optimizing the difference between the prediction of clean example $f(x)$ and that of adversarial example $f^{k}(x^{\prime})$. And there is a hyper-parameter balancing the importance of the two parts.
@@ IV-A Evaluation Setting
CIFAR-10 Wu, Xia, and Wang[wu2020adversarial] WRN-28-10(RST-AWP) 88.3% 60.0%
Gowal et al.[gowal2020uncovering] WRN-70-16 85.3% 57.2%
Rebuffi et al.[rebuffi2021data] WRN-28-10 86.1% 57.5%
Rebuffi et al.[rebuffi2021data] WRN-34-10 86.2% 58.1%
Rebuffi et al.[rebuffi2021data] WRN-70-16 87.3% 60.1%
Sehwag et al.[sehwag2021robust] + 10M DDPM WRN-34-10 86.7% 60.3%
Pang et al.[pang2022robustness] + 1M DDPM WRN-28-10 88.6% 61.0%
Pang et al.[pang2022robustness] + 1M DDPM WRN-70-16 89.0% 63.4%
Jia et al.[jia2022adversarial] WRN-34-10(LAS-AWP) 87.7% 55.5%
Jia et al.[jia2022adversarial] WRN-70-16 85.7% 57.6%
@@
CIFAR-100 Wu, Xia, and Wang[wu2020adversarial] WRN-34-10 60.4% 28.9%
Gowal et al.[gowal2020uncovering] WRN-70-16 60.9% 30.0%
Rebuffi et al.[rebuffi2021data] WRN-28-10 63.0% 29.8%
Rebuffi et al.[rebuffi2021data] WRN-70-16 65.8% 32.4%
Sehwag et al.[sehwag2021robust] + 1M DDPM WRN-34-10 65.9% 31.2%
Pang et al.[pang2022robustness] + 1M DDPM WRN-28-10 63.7% 31.1%
Pang et al.[pang2022robustness] + 1M DDPM WRN-70-16 65.6% 33.1%
Jia et al.[jia2022adversarial] WRN-34-10 64.9% 30.8%
@@
SVHN Wu, Xia, and Wang[wu2020adversarial] WRN-34-10 92.8% 58.7%
Gowal et al.[gowal2020uncovering] WRN-28-10 92.9% 57.3%
Gowal et al.[gowal2020uncovering] WRN-34-10 93.1% 58.0%
Rebuffi et al.[rebuffi2021data] WRN-28-10 94.5% 57.3%
@@
FGSM, PGD, and AutoAttack are utilized to evaluate our proposed method along with state-of-the-arts. Testing robustness against PGD is done in 40/100 iterations with the maximum $\epsilon=8/255$ and step sizes of $0.01$ for each iteration. FGSM attack also uses the same 8/255 bound for perturbation. We leverage the implementation of AutoAttack from torchattacks repository [kim2020torchattacks].
Model training details. We use WideResNets as our base models: WRN-28-10, WRN-34-10 and WRN-70-16 [zagoruyko2016wide], respectively. Each branch is trained for 50 epochs using stochastic gradient descent (SGD) as the optimizer, with an initial learning rate of 0.1 that decays to 0.01 at the 40th epoch. The hyper-parameters are configured as $\lambda_{1}=\lambda_{2}=3$ during the first 40 epochs, followed by $\lambda_{1}=\lambda_{2}=1$ for the remaining 10 epochs.
@@
TABLE III: Ablation studies of loss terms. Vanilla: Models trained entirely with clean samples. TRADES-4: The same model structure as our method proposed, but the branches are trained according to TRADES[zhang2019theoretically](differ in loss function).
@@ IV-B Overall Performance
To illustrate the effectiveness of the proposed BORT, we train various models using this framework and evaluate their robustness against different adversarial attack algorithms. Furthermore, the proposed method is compared with different state-of-the-art approaches [wu2020adversarial, gowal2020uncovering, jia2022adversarial, sehwag2021robust, pang2022robustness, rebuffi2021data]. The detailed experiment results are illustrated in Table I. For a fair comparison, only the results that do not use the additional dataset (80M-Ti for example) are presented in the table, although we still outperform the results using it. The 80M-Ti dataset has been taken down by the author team, given that it may contain offensive and biased content. Likewise, we do not try to train stronger models utilizing this dataset as additional training data.
@@ V Conclusion
In this paper, we present BORT, the first adversarial training framework leveraging a multi-branch neural network, which improves network representation through multiple orthogonal branches while keeping inference time and computational cost from increasing, greatly improving model robustness against adversarial attacks.